OpenStudy (anonymous):

what is the integration of sec2x.cosx

OpenStudy (anonymous):

\[\sec^2x=\frac{1}{\cos^2x}\] So \[\sec^2x*\cos x=\frac{1}{\cos^2x}*\cos x=\frac{\cos x}{\cos^2x}=\frac{1}{\cos x}=\sec x\]

OpenStudy (anonymous):

Next step would be to integrate sec x \[ \int \sec x dx \] Ok using the table in the back of a calc book we see that the \[ \int \sec x dx = \ln | \sec x + \tan x | +C\]